### v1

@@ 4 Discussion
That said, our ability to definitively trace the mechanisms behind peer influence is limited by the nature of our historical data. Our specification that examines cadets’ choices based on the composition of their cohort by the states of origin of their peers enables us to causally identify peer influence. Nonetheless, it does not provide insight into why peers influence cadets’ decisions. One curious aspect is that the peer influence only affected cadets from states with slaves. This could be due to the fact that West Point was an institution in which cadets were being trained for the U.S. Army, and so only cadets from Slave States faced a conflicting allegiance. Our evidence on how joint participation in the Mexican-American War amplifies peer influence suggests that sharing experiences matters. Theoretically, however, we are still left with the fact that peer effects could have operated in at least three ways: (1) by fostering communication and persuasion, whereby cadets from Slave States may have been convinced to support the Union cause through exposure to pro-Union peers, (2) by building friendship bonds that influenced cadets’ decisions to fight alongside their peers, or (3) by shifting perceptions of the likely outcome of the war, where cadets from Free States could have led their peers to believe that the Union was more likely to prevail. Our data, however, do not allow us to disentangle these, or other mechanisms, and all could have operated together (which would be consistent with some evidence from personal letters and diaries (McPherson, 1997)). Further investigation of peer influence on allegiances in modern contexts, where additional data are more readily available, could offer additional insights into how such mechanisms operate.
Data Availability: all data will be made available upon publication.
Code Availability: all code will be made available upon publication.
@@ A.2 Classification of States
South Carolina 57.2 Slave-State $>$5% $>$10% Heavy-Slave State
Mississippi 55.2 Slave-State $>$5% $>$10% Heavy-Slave State
Louisiana 46.9 Slave-State $>$5% $>$10% Heavy-Slave State
Alabama 45.1 Slave-State $>$5% $>$10% Heavy-Slave State
Florida 44.0 Slave-State $>$5% $>$10% Heavy-Slave State
Georgia 43.7 Slave-State $>$5% $>$10% Heavy-Slave State
North Carolina 33.7 Slave-State $>$5% $>$10% Heavy-Slave State
Virginia 30.7 Slave-State $>$5% $>$10% Mid-Slave State
Texas 30.2 Slave-State $>$5% $>$10% Mid-Slave State
Arkansas 25.5 Slave-State $>$5% $>$10% Mid-Slave State
Tennessee 24.8 Slave-State $>$5% $>$10% Mid-Slave State
Kentucky 19.5 Slave-State $>$5% $>$10% Border-State
Maryland 12.7 Slave-State $>$5% $>$10% Border-State
Missouri 9.7 Slave-State $>$5% $<$10% Border-State
District of Columbia 4.4 Slave-State $<$5% $<$10% Border-State
Delaware 1.6 Slave-State $<$5% $<$10% Border-State
New Jersey 0.01 Free-State $<$5% $<$10% Free-State
New York 0.0 Free-State $<$5% $<$10% Free-State
Pennsylvania 0.0 Free-State $<$5% $<$10% Free-State
Ohio 0.0 Free-State $<$5% $<$10% Free-State
Illinois 0.0 Free-State $<$5% $<$10% Free-State
Indiana 0.0 Free-State $<$5% $<$10% Free-State
Massachusetts 0.0 Free-State $<$5% $<$10% Free-State
Wisconsin 0.0 Free-State $<$5% $<$10% Free-State
Michigan 0.0 Free-State $<$5% $<$10% Free-State
Iowa 0.0 Free-State $<$5% $<$10% Free-State
Maine 0.0 Free-State $<$5% $<$10% Free-State
Connecticut 0.0 Free-State $<$5% $<$10% Free-State
California 0.0 Free-State $<$5% $<$10% Free-State
New Hampshire 0.0 Free-State $<$5% $<$10% Free-State
Vermont 0.0 Free-State $<$5% $<$10% Free-State
Rhode Island 0.0 Free-State $<$5% $<$10% Free-State
Minnesota 0.0 Free-State $<$5% $<$10% Free-State
Oregon 0.0 Free-State $<$5% $<$10% Free-State
@@ A.3 Summary Statistics
Panel A: Heavy-Slave States (slave share $>$ 33%)
@@
Panel B: Border and Mid-Slave States ( 1%$<$slave share $\leq$ 33%)
@@
Panel C: Free States (slave share $<$ 1%)
@@ Appendix B Additional Results
B.1 Alternative Thresholds
@@
Share Free-State peers (sd) 0.052∗∗ 0.053∗∗ -$0$0.004 0.059∗∗ 0.058∗∗ -0.002 -0.005
@@
B.2 Home State Economic Proxies and War Allegiances
@@
B.3 Peer influence and shared Mexican-American War experience
@@
B.4 Peer Influence by Cohort
Figure S3: Peer influence by cohort
B.5 Peer Influence pre and post 1850
@@
B.6 Graduate Peers vs. Dropout Peers
@@
B.7 Multinomial Logit Results
@@
B.8 Considering Cash-Crop Employment
@@
Share Free-State peers (sd) 0.062∗∗ 0.044∗∗ 0.039 0.006 -$0$0.003
@@
Share Free-State peers (sd) 0.042 0.037 -$0$0.003
@@
B.9 Considering Appointment State Peers
@@
B.10 Correlation between Academic Rank and Military Rank in 1865
Dependent var Rank in 1865: war participants
The Confederacy The Union
(1) (2) (3) (4) (5) (6)
Class rank $0$0.001 -0.001 0.006∗∗∗ 0.006∗∗∗ 0.006∗∗
(0.003) (0.003) (0.003) (0.002) (0.002) (0.003)
Age in 1860 -0.004 -0.017 -0.034 -0.039
(0.063) (0.068) (0.043) (0.045)
Slave pop. share (sd) 0.148 0.029
(0.089) (0.063)
Cohort -0.049 -0.065 -0.087∗∗ -0.092∗∗
(0.061) (0.066) (0.042) (0.044)
State FEs N N Y N N Y
Dependent var.mean 7.516 7.516 7.516 5.848 5.848 5.848
Observations 281 281 281 637 637 637
R-squared 0.000 0.085 0.156 0.010 0.108 0.127
Note. The sample for this analysis consists of individuals who served in the Confederate Army (Columns (1)-(3)) and those who served in the Union Army (Columns (4)-(6)). This table shows that the association between academic rank and military rank is clear in the Union but not in the Confederacy. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.

### v2

@@ 4 Discussion
That said, our ability to definitively trace the mechanisms behind peer influence is limited by the nature of our historical data. Our specification that examines cadets’ choices based on the composition of their cohort by the states of origin of their peers enables us to causally identify peer influence. Nonetheless, it does not provide insight into why peers influence cadets’ decisions. One curious aspect is that the peer influence only affected cadets from states with slaves. This could be due to the fact that West Point was an institution in which cadets were being trained for the U.S. Army, and so only cadets from Slave States faced a conflicting allegiance. Our evidence on the strength of within-cohort influence and how joint participation in the Mexican-American War amplifies peer influence suggests that shared-experiences mattered.
Theoretically, however, we are still left with the fact that peer effects could have operated in at least three ways: (1) by fostering communication and persuasion, whereby cadets from Slave States may have been convinced to support the Union cause through exposure to pro-Union peers, (2) by building friendship bonds that influenced cadets’ decisions to fight alongside their peers, or (3) by shifting perceptions of the likely outcome of the war, where cadets from Free States could have led their peers to believe that the Union was more likely to prevail. Our data, however, do not allow us to disentangle these, or other mechanisms, and all could have operated together (which would be consistent with some evidence from personal letters and diaries (McPherson, 1997)). Further investigation of peer influence on allegiances in modern contexts, where additional data are more readily available, could offer additional insights into how such mechanisms operate.
Data and Code Availability: all data and code will be made available upon publication.
@@ A.1 Data Construction
Participation in the Mexican-American War
The information is constructed from the Biographical Register of the Officers and Graduates of the United States Military Academy. We classify as participants those who were directly engaged in combat or military operations during the Mexican-American War, as well as those who are listed as contributing indirectly to the war effort—for instance, by providing logistical support such as transporting supplies.
@@ A.2 Classification of States
South Carolina 57.2 Slave-State >>5% >>10% Heavy-Slave State
Mississippi 55.2 Slave-State >>5% >>10% Heavy-Slave State
Louisiana 46.9 Slave-State >>5% >>10% Heavy-Slave State
Alabama 45.1 Slave-State >>5% >>10% Heavy-Slave State
Florida 44.0 Slave-State >>5% >>10% Heavy-Slave State
Georgia 43.7 Slave-State >>5% >>10% Heavy-Slave State
North Carolina 33.7 Slave-State >>5% >>10% Heavy-Slave State
Virginia 30.7 Slave-State >>5% >>10% Mid-Slave State
Texas 30.2 Slave-State >>5% >>10% Mid-Slave State
Arkansas 25.5 Slave-State >>5% >>10% Mid-Slave State
Tennessee 24.8 Slave-State >>5% >>10% Mid-Slave State
Kentucky 19.5 Slave-State >>5% >>10% Border-State
Maryland 12.7 Slave-State >>5% >>10% Border-State
Missouri 9.7 Slave-State >>5% <<10% Border-State
District of Columbia 4.4 Slave-State <<5% <<10% Border-State
Delaware 1.6 Slave-State <<5% <<10% Border-State
New Jersey 0.01 Free-State <<5% <<10% Free-State
New York 0.0 Free-State <<5% <<10% Free-State
Pennsylvania 0.0 Free-State <<5% <<10% Free-State
Ohio 0.0 Free-State <<5% <<10% Free-State
Illinois 0.0 Free-State <<5% <<10% Free-State
Indiana 0.0 Free-State <<5% <<10% Free-State
Massachusetts 0.0 Free-State <<5% <<10% Free-State
Wisconsin 0.0 Free-State <<5% <<10% Free-State
Michigan 0.0 Free-State <<5% <<10% Free-State
Iowa 0.0 Free-State <<5% <<10% Free-State
Maine 0.0 Free-State <<5% <<10% Free-State
Connecticut 0.0 Free-State <<5% <<10% Free-State
California 0.0 Free-State <<5% <<10% Free-State
New Hampshire 0.0 Free-State <<5% <<10% Free-State
Vermont 0.0 Free-State <<5% <<10% Free-State
Rhode Island 0.0 Free-State <<5% <<10% Free-State
Minnesota 0.0 Free-State <<5% <<10% Free-State
Oregon 0.0 Free-State <<5% <<10% Free-State
@@ A.3 Summary Statistics
Panel A: Heavy-Slave States (slave share >> 33%)
@@
Panel B: Border and Mid-Slave States ( 1%<<slave share $\leq$ 33%)
@@
Panel C: Free States (slave share << 1%)
@@
Table S3: Summary Statistics: Cohort Level
Mean Std. Dev. Min. Max. Obs.
Panel A: Heavy-Slave States (slave share >> 33%)
Number of graduates 5.146 2.516 0 11 41
Fraction of graduates 0.127 0.056 0 0.244 41
Panel B: Border and Mid-Slave States ( 1%<<slave share $\leq$ 33%)
Number of graduates 11.293 3.907 3 19 41
Fraction of graduates 0.282 0.080 0.088 0.480 41
Panel C: Free States (slave share << 1%)
Number of graduates 23.512 5.459 12 33 41
Fraction of graduates 0.591 0.084 0.462 0.794 41
Note. This table presents summary statistics at the cohort level.
@@ Appendix B Additional Results
B.1 Cohort-Clustered and Bootstrap Cohort-Clustered SE
Table S5: Peer composition and allegiance choice: variations on standard errors
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4) (5) (6)
A. Cohort level bootstrap
Share Free-State peers (sd) 0.062∗∗∗ 0.045∗∗∗ 0.040∗∗ 0.006 0.004 0.002
(0.022) (0.017) (0.017) (0.015) (0.014) (0.015)
Contronls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.220 0.268 0.000 0.015 0.028
B. Cohort level cluster
Share Free-State peers (sd) 0.062∗∗∗ 0.045∗∗∗ 0.040∗∗ 0.006 0.004 0.002
(0.022) (0.016) (0.020) (0.012) (0.012) (0.012)
Contronls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.017 0.220 0.268 0.000 0.015 0.028
Note. Panel A reports the re-estimation of the baseline results using robust standard errors clustered at the cohort level. Panel B reports the corresponding re-estimates using bootstrapped standard errors clustered at the cohort level with 400 resampling iterations. (This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. ***$p<0.01$, **$p<0.05$, *$p<0.1$.)
B.2 Alternative Thresholds
@@
Share Free-State peers (sd) 0.052∗∗ 0.053∗∗ -0.000 -0.004 0.059∗∗ 0.058∗∗ -0.002 -0.005
@@
B.3 Home State Economic Proxies and War Allegiances
@@
B.4 Correlation between Academic Rank and Military Rank in 1865
Dependent var Rank in 1865: war participants
The Confederacy The Union
(1) (2) (3) (4) (5) (6)
Class rank 0.000 -0.001 -0.001 0.006∗∗∗ 0.006∗∗∗ 0.006∗∗
(0.003) (0.003) (0.003) (0.002) (0.002) (0.003)
Age in 1860 -0.004 -0.017 -0.034 -0.039
(0.063) (0.068) (0.043) (0.045)
Slave pop. share (sd) 0.148 0.029
(0.089) (0.063)
Cohort -0.049 -0.065 -0.087∗∗ -0.092∗∗
(0.061) (0.066) (0.042) (0.044)
State FEs N N Y N N Y
Dependent var.mean 7.516 7.516 7.516 5.848 5.848 5.848
Observations 281 281 281 637 637 637
R-squared 0.000 0.085 0.156 0.010 0.108 0.127
Note. The sample for this analysis consists of individuals who served in the Confederate Army (Columns (1)-(3)) and those who served in the Union Army (Columns (4)-(6)). This table shows that the association between academic rank and military rank is significant in the Union but not in the Confederacy. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.5 Peer Influence by Cohort
Figure S3: Peer influence by cohort
B.6 Peer influence and shared Mexican-American War experience
@@
B.7 Peer Influence pre and post 1850
@@
B.8 Controlling for Five-Year Fixed Effects
Table S11: Peer Composition and Allegiance Choice with Five-year Fixed Effects
Dependent var Join the Union: war participants
Slave States Free States
(1) (2) (3) (4) (5) (6)
Share Free-State peers (sd) 0.061∗∗ 0.048∗∗ 0.043 0.002 0.002 -0.001
(0.026) (0.023) (0.024) (0.013) (0.013) (0.013)
Contronls N Y Y N Y Y
State FEs N N Y N N Y
Five-Year Interval FEs Y Y Y Y Y Y
Dependent var. mean 0.362 0.362 0.362 0.919 0.919 0.919
Observations 389 389 389 540 540 540
R-squared 0.031 0.225 0.272 0.033 0.041 0.055
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. We include fixed effects at five year intervals to allow for a shift in the environment. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
B.9 Graduate Peers vs. Dropout Peers
@@ B.8 Controlling for Five-Year Fixed Effects
B.10 Multinomial Logit Results
@@
B.11 Considering Cash-Crop Employment
@@
Share Free-State peers (sd) 0.062∗∗ 0.044∗∗ 0.039 0.006 -0.000 -0.003
@@
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Following White (2024), we introduce a variable Cash crops, defined as a dummy equal to 1 if the cadet had a history of employment in cash-crop agriculture. This includes cases where the graduate was recorded as having spent time as a planter (plantation owner) or held any occupation related to the production, processing, sale, or export of cotton, indigo, rice, sugar, or tobacco. Columns (1)-(3) focus on cadets from Slave States and Columns (4)-(6) on those from Free States. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@
Share Free-State peers (sd) 0.042 0.037 -0.000 -0.003
@@
B.12 Considering Appointment State Peers